OpenStudy (anonymous):

simplify (x²y³z)/(4x) divided by (3xy⁴ z²)/(2z)

OpenStudy (jdoe0001):

\(\bf \cfrac{ \frac{x^2y^3z}{4x} }{ \frac{3xy^4z^2}{2z} }\qquad \qquad recall\implies \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c} \\ \quad \\ \cfrac{ \frac{x^2y^3z}{4x} }{ \frac{3xy^4z^2}{2z} }\implies \cfrac{x^2y^3z}{4x}\cdot \cfrac{2z}{3xy^4z^2}\implies \cfrac{2zx^2y^3z}{4x\cdot 3xy^4z^2}\)